OpenStudy (anonymous):

Help please solve........ -15(4+3x)-8x=26-2(x-9)

OpenStudy (saifoo.khan):

-60 -45x -8x = 26 -2x + 19 Can u solve now?

OpenStudy (anonymous):

x=-2.05

OpenStudy (anonymous):

its -60 -45x-8x=26-2x+18

OpenStudy (anonymous):

sox\[\approx\]-2.04

OpenStudy (anonymous):

Thank you